OpenStudy (anonymous):

What is the standard form of a line that goes through the points 3,-6 and -2.-1 ?

OpenStudy (anonymous):

you need the sloe first. slope is \[m=\frac{y_2-y_2}{x_2-x_1}=\frac{-1+6}{-2-3}=-1\]

OpenStudy (anonymous):

The standard equation of a line is y=mx + b Where m is the slope of the line and b is the y intercept of the line To find slope we use (y2-y1) / x2-x1) which gives us a slope of 5/-5 or -1 m = -1 To find b we substitute the m we found into the equation and use one of the points given. -6 = -1(3) +b -6 = -3 + b -3 = b Therefore the standard equation of a line for this case is y= -x -3

OpenStudy (anonymous):

then use point-slope formula \[y-y_1=m(x-x_1)\] with \[m=-1, x_1=-2, y_1=-1\] to get \[y+1=-(x+2)\] multiply on the right to get \[y+1=-x+2\]

OpenStudy (anonymous):

unfortunately i think "standard form" is \[ax+by+c=0\] maybe

OpenStudy (anonymous):

so write \[x+y-1=0\] but of course it is all the same

OpenStudy (anonymous):

\[y=mx+b\] is "slope- intercept" form \[ax+by+c=0\] is "standard form" and \[y-y_1=m(x-x_1)\] is "point-slope" form

OpenStudy (anonymous):

although i certainly prefer \[y=mx+b\] because it is easy to compute points if you just pick a number for x and plug it in